OpenStudy (anonymous):

1.Find the slope for the equation -2x + 3y = 12

OpenStudy (nincompoop):

to find the slope, first convert the equation into slope-intercept form y = mx + b whereby m would be your slope

OpenStudy (nincompoop):

so for example, you are given an equation of 6x - 9y = 36 isolate y on one side by subtracting 6x on both sides of the equation 6x - 6x - 9y = 36 - 6x -9y = -6x +36 since we want just the y we are going to divide both sides by -9 y = 1/3 x + 4 then the slope would be 1/3

OpenStudy (anonymous):

so I would subtract -2x from both sides then I divide 3y

OpenStudy (nincompoop):

correct, subtracting -2x would be adding 2x since - (-2x) is +2x

OpenStudy (nincompoop):

the technique is that whatever you're trying to get rid on the one side, use the opposite of it so if you have a minus, then you add, if you have a multiplication, then divide... and so on

OpenStudy (anonymous):

ok so would be the equation change to 3y = 2x + 12

OpenStudy (nincompoop):

correct, but you need just y not 3y in order to achieve that you will have to do what? and remember that whatever you do one one side, must be done on the other side as well

OpenStudy (anonymous):

ok, the dividing part is kinda hard for me on this one could you help me out

OpenStudy (anonymous):

@nincompoop

OpenStudy (anonymous):

would it be y = 2/3x + 4 @nincompoop

OpenStudy (nincompoop):

correct